# A Flat Minima Perspective on Understanding Augmentations and Model Robustness

### v1

@@ 4.3 Tests on Model Flatness
Dataset Metrics      ERM      AugMix      RandAug      PixMix      DeepAug      StyleAug CIFAR-10 $\mu_{\text{PAC-Bayes}}\downarrow$ 166.72 116.70 (-50.02) 113.68 (-53.04) 105.30 (-61.42) 66.89 (-99.83) 122.76 (-43.96) LPF $\downarrow$ 1.01 0.52 (-0.49) 0.46 (-0.55) 0.37 (-0.64) 0.71 (-0.30) 1.14 (+0.13) $\epsilon_{\text{sharp}}$downarrow 10.28 7.42 (-2.86) 7.32 (-2.96) 9.19 (-1.09) 11.13 (+0.85) 19.84 (+9.56) CIFAR-100 $\mu_{\text{PAC-Bayes}}\downarrow$ 179.92 155.74 (-24.18) 151.15 (-28.77) 136.07 (-43.85) 71.56 (-108.36) 135.48 (-44.44) LPF $\downarrow$ 2.37 1.87 (-0.50) 1.67 (-0.70) 1.40 (-0.97) 2.38 (+0.01) 3.39 (+1.02) $\epsilon_{\text{sharp}}$downarrow 8.98 8.35 (-0.63) 7.95 (-1.03) 8.00 (-0.98) 13.49 (+4.51) 42.82 (+33.84)
4.4 Tests on Generalization Capability
@@
We finally investigate whether augmentations having rich proximal density indeed fosters model robustness against diverse distribution shifts (Theorem 3).
Benchmarks to be considered. Generalization refers to a model’s ability to perform well on unseen data, which here spans natural corruptions of clean samples, adversarial attacks, and distributional changes from domain shifts. Given that prior work has already explored simpler train-test generalization scenarios (with limited distributional discrepancies), we focus on more challenging shifts, such as common corruptions and adversarial attacks on CIFAR CIFAR , tinyImageNet, and ImageNet IN-C7:2019 , as well as domain generalization using PACS PACS and OfficeHome OfficeHome .
Robustness Benchmarks and Metrics. Following the setup in Table 2, we train our models on CIFAR-10/100, tinyImageNet, and ImageNet datasets and then evaluate their robustness to common corruptions and untargeted PGD $L_{2}/L_{\infty}$ adversarial attacks. For domain generalization, we adopt the approach of DomainBed DomainBed , employing ResNet50 and reporting average accuracies over three independent runs for each benchmark and augmentation strategy. In Table 3, lower values represent better performance on corruption and adversarial benchmarks, while higher values indicate better domain generalization. See Appendix C.1.2 for further details on benchmarks, metrics, and training.
Experimental Results. The outcomes align closely with Theorems 1 to 3. Specifically, every augmentation method that maintained nonzero probability near the original image surpasses ERM in overall robustness. Moreover, the more densely each method populate the proximity region, the larger the performance boost. For instance, AugMix populates many augmented samples near the originals in both CIFAR and tinyImageNet, which translates into top-level robustness. RandAug and PixMix keep only a few proximal variants, enough to outperform ERM but not to dominate. DeepAug leaves proximal region empty in CIFAR, yet fills the region succesfully on ImageNet. Therfore, DeepAug underperforms on CIFAR while excelling on (tiny)ImageNet. StyleAug produces no nearby samples at all and yields no robustness gains.
We interpret these observations as the evidence of the relationship between augmentations and robustness strongly holding, when the nonzero probability condition (Assumption 1) is met.The results also co-aligns with our theorems, highlighting the importance of rich proximal density in improving model generalization capability.
@@
Table 3: Comparison of various augmentation methods against ERM across multiple robustness scenarios, including common corruptions (CIFAR-10/100-C/$\overline{\text{C}}$, tinyImageNet-C, ImageNet-C), adversarial attacks ($L_{2}$, $L_{\infty}$), and domain shifts (PACS, OfficeHome). Metrics with ()($\downarrow$)( ↓ ) favor lower values (corruptions and adversarial) and ()($\uparrow$)( ↑ ) favor higher values (domain generalization). Differences from ERM are shown in parentheses.
Robustness Benchmark ERM      AugMix      RandAug      PixMix      DeepAug      StyleAug CIFAR-10-C $\downarrow$ 30.54 15.24 (-15.30) 19.65 (-10.89) 10.57 (-19.97) 31.92 (+1.38) 30.50 (-0.04) CIFAR-10-$\overline{\text{C}}$downarrow 8.13 1.73 (-6.40) 3.07 (-5.06) 1.05 (-7.08) 17.99 (+9.86) 9.34 (+1.21) Common CIFAR-100-C $\downarrow$ 59.05 42.64 (-16.41) 46.59 (-12.46) 35.14 (-23.91) 62.34 (+3.29) 70.92 (+11.87) Corruption CIFAR-100-$\overline{\text{C}}$downarrow 18.32 5.01 (-13.31) 8.31 (-10.01) 4.78 (-13.54) 39.54 (+21.22) 30.77 (+12.45) tinyImageNet-C $\downarrow$ 74.58 63.47 (-11.11) 66.95 (-7.63) 61.51 (-13.07) 62.51 (-12.07) 79.31 (+4.73) ImageNet-C $\downarrow$ 69.37 69.34 (-0.03) 66.88 (-2.49) 61.51 (-7.86) 56.68 (-12.69) 75.52 (+6.15) CIFAR-10, $L_{2}$downarrow 77.67 70.68 (-6.99) 76.30 (-1.37) 65.24 (-12.43) 91.07 (+13.40) 91.13 (+13.46) CIFAR-10, $L_{\infty}$downarrow 98.48 99.57 (+1.09) 99.86 (+1.38) 99.46 (+0.98) 100.00 (+1.52) 99.97 (+1.49) CIFAR-100, $L_{2}$downarrow 98.72 92.78 (-5.94) 95.92 (-2.60) 90.11 (-8.61) 98.37 (-0.35) 99.67 (+0.95) Adversarial CIFAR-100, $L_{\infty}$downarrow 99.94 99.67 (-0.27) 99.93 (-0.01) 99.60 (-0.34) 99.99 (+0.05) 99.98 (+0.04) Attacks tinyImageNet $L_{2}$downarrow 57.89 61.12 (+3.23) 58.86 (+0.97) 63.57 (+5.68) 52.91 (-4.98) 83.50 (+25.61) tinyImageNet $L_{\infty}$downarrow 99.91 99.89 (-0.02) 99.95 (+0.04) 99.94 (+0.03) 99.92 (+0.01) 100.00 (+0.09) ImageNet, $L_{2}$downarrow 76.16 76.06 (-0.10) 76.09 (-0.07) 75.44 (-0.72) 75.37 (-0.79) 87.71 (+11.55) ImageNet, $L_{\infty}$downarrow 99.70 99.69 (-0.01) 99.70 (+0.00) 99.63 (-0.07) 99.62 (-0.08) 99.92 (+0.22) Domain PACS $\uparrow$ 79.8 82.1 (+2.3) 83.4 (+3.6) 79.5 (-0.3) 64.4 (-15.4) 83.7 (+3.9) Shifts OfficeHome $\uparrow$ 68.2 68.5 (+0.3) 67.4 (-0.8) 63.7 (-4.5) 44.7 (-23.5) 64.5 (-3.7)
5 Discussion and Future Work
@@
  • We first offer a flat minima perspective to elucidate how augmentations affect robustness.
@@
    • A weak duality between adversarial training (AT) and sharpness-aware minimization (SAM) is known to exist zhang2024duality . Nonetheless, the result holds only for linear models with Gaussian data and fails to cover standard augmentations.
@@
  • We first establish a formal connection between the general form of data augmentations and enhanced robustness under general distribution shifts.
@@
    • Previous studies faced limitations such as focusing solely on: i) experimental analysis, ii) specific augmentations, or iii) certain distributional shifts.
@@
    • In contrast, our analysis imposes no such restrictions on augmentation types or benchmarks. We validate our findings through experiments across diverse robustness benchmarks.
Despite these contributions, our work has certain limitations. Specifically, we have not conducted massive experiments to validate our theorem across a broader range of benchmarks and augmentation methods. For instance:
  • Our theoretical framework does not encompass label-manipulating augmentations like Mixup.
  • Sophisticated adversarial attacks such as the C&W attack remain unexplored in our experiments.
Future research may advance our work by incorporating label-manipulating augmentations and examining a more extensive set of techniques. For further discussion of theorem applications, flat-minima optimizers, and adversarial training, see Appendix E, which also provides interpretation to augmentation compositions based on our findings.
6 Related Works
In this section, we provide a brief overview of previous works on data augmentation and model robustness. For more detailed discussions on prior works, refer to Appendix D.
Augmentation Strategies. Augmentations can be grouped as model-free, model-based, or policy-based augmentations augment_survey . Model-free methods act on one image (translation, rotation, masking CutOut ; HideNSeek ) or several images (Mixup, CutMix Mixup:2018 ; CutMix:2019 ). Advanced augmentations (AugMix, PixMix AugMix9:2021 ; PixMix:2022 ) consisting of multiple image augmentation steps, are known to improve corruption robustness AugMix9:2021 ; PixMix:2022 . Model-based approaches rely on pretrained generators (CGAN and its variants CGAN ; CGAN_var1 ; CGAN_var2 ; CGAN_var3 ; CGAN_var4 ) or image-to-image networks (DeepAugment, ANT, StyleAug) to generate augmented images. Policy-based schemes learn augmentation policies: AutoAugment AutoAugment and its successors FastAA ; FasterAA ; RandAugment ; AdaTransform ; AdvAA ; AugMax fall into this group. Despite such augmentations’ strong empirical gains, a unifying theory on understanding their impacts on robustness still missing.
Augmentation and Robustness. A number of previous studies tries to find the link between augmentation and robustness in several angles. A game-theoretic view exists on interpreting augmentation’s effect on common corruption robustness aug_analysis_cc1 . A Fourier perspective on defining adversarial and common corruption robustness exists AugFourier . Interpretations on adversarial perturbation as the worst-case distribution shifts in the data ME-ADA ; ADA exist. Certain augmentations with weight averaging is known to improve adversarial and common corruption robustness AugmentNAdvGood ; boundary_thickness . Yet, these explanations either remain empirical or tied to specific augmentations, lacking applicability in general settings.
Beyond Augmentations. The duality between adversarial training and sharpness-aware minimization optimizer is known to exist on a linear classifier over Gaussian distributions zhang2024duality . Preserving fairness under distribution shifts by mimicking the shift as perturbations utilizing transportation function has been studied jiang2023chasing . Flatter minima is shown to mitigate boundary tilting problem boundary_tilting and is empirically found to be strong against common corruptions flat_minima_cc . Still, none of these works can explain augmentation-driven robustness.
7 Conclusion
Our framework links data augmentation to robustness through the lens of flat minima under general distribution shifts. When augmentations densely fill the neighborhood of an image, they map to broad, flat regions in parameter space, tightening generalization bounds and improving robustness to distribution shifts. Sparse augmentations, by contrast, leave gaps near the original image and confer little protection. Experiments on CIFAR, ImageNet, PACS, and OfficeHome—covering common corruptions, adversarial attacks, and domain shifts—corroborate our theory. We hope these insights guide the design of next-generation augmentation methods, including those built on foundation models.
8 Acknowledgment

### v3

@@ Introduction
Dataset Metrics     ERM    AugMix    RandAug    PixMix    DeepAug    StyleAug CIFAR-10 $\mu_{\text{PAC-Bayes}}\downarrow$ 168.83 117.16 (-51.67) 110.80 (-58.03) 102.87 (-65.96) 67.74 (-101.09) 122.80 (-46.03) LPF $\downarrow$ 1.43 0.54 (-0.89) 0.45 (-0.98) 0.37 (-1.06) 0.72 (-0.71) 1.13 (-0.30) $\epsilon_{\text{sharp}}\downarrow$ 40.90 25.16 (-15.74) 24.47 (-16.43) 26.57 (-14.33) 55.63 (+14.73) 130.64 (+89.74) CIFAR-100 $\mu_{\text{PAC-Bayes}}\downarrow$ 181.10 155.74 (-25.36) 149.79 (-31.31) 134.58 (-46.52) 72.24 (-108.86) 129.02 (-52.08) LPF $\downarrow$ 2.37 1.84 (-0.53) 1.69 (-0.68) 1.42 (-0.95) 2.37 (0.00) 3.95 (+1.58) $\epsilon_{\text{sharp}}\downarrow$ 42.39 39.32 (-3.07) 37.76 (-4.63) 33.14 (-9.25) 84.48 (+42.09) 302.81 (+260.42) Improvement Rate 6/6 6/6 6/6 3/6 2/6
Table 2: Flatness metrics of ERM and different augmentations on CIFAR datasets. Overall, when augmentation meets the PSA condition (nonzero probability near the original image) strongly, the flatter minima are found consistently. ($\downarrow$: The lower the better, i.e., flatter minimum.)
   Benchmarks ERM    AugMix    RandAug    PixMix    DeepAug    StyleAug CIFAR-10-C $\downarrow$ 30.54 15.24 (-15.30) 19.65 (-10.89) 10.60 (-19.94) 31.92 (+1.38) 30.50 (-0.04) CIFAR-10-$\overline{\text{C}}$ $\downarrow$ 31.35 20.28 (-11.07) 20.64 (-10.71) 14.60 (-16.75) 36.75 (+5.40) 36.57 (+5.22) CIFAR-100-C $\downarrow$ 59.04 42.64 (-16.40) 46.59 (-12.45) 35.20 (-23.84) 62.34 (+3.30) 70.91 (+11.87) CIFAR-100-$\overline{\text{C}}$ $\downarrow$ 62.43 48.38 (-14.05) 48.32 (-14.11) 40.20 (-22.23) 67.91 (+5.48) 76.56 (+14.13) CIFAR-10, $L_{2}$ $\downarrow$ 77.61 70.76 (-6.85) 76.15 (-1.46) 65.81 (-11.80) 91.18 (+13.57) 91.06 (+13.45) CIFAR-10, $L_{\infty}$ $\downarrow$ 98.49 99.10 (+0.61) 99.86 (+1.37) 99.46 (+0.97) 100.00 (+1.51) 99.98 (+1.49) CIFAR-100, $L_{2}$ $\downarrow$ 98.73 92.76 (-5.97) 96.06 (-2.67) 90.69 (-8.04) 98.44 (-0.29) 99.69 (+0.96) CIFAR-100, $L_{\infty}$ $\downarrow$ 99.94 99.67 (-0.27) 99.94 (0.00) 99.69 (-0.25) 99.99 (+0.05) 99.98 (+0.04)     Improvement Rate 7/8 6/8 6/8 1/8 1/8
[…]
Empirical Examination on Proximal Density, Model Flatness, and Robustness
To empirically validate our theory, we examine proximal density, model flatness, and robustness on different benchmarks. Earlier, we established that any label-invariant augmentation assigning nonzero probability density to the neighborhood of the original image (PSA condition) maps to an analogous region in parameter space (Theorem 1), which in turn induces flatter minima (Theorem 2) and yields a tighter generalization bound (Theorem 3). Therefore, when augmentation populates rich proximal representations for a given dataset (PSA condition strongly met), it should yield flat minima and strong robustness gains under distribution shifts. In contrast, when augmentation fail to bring robustness, the PSA condition will not be satisfied.
Augmentation Methods to Be Considered
Following the taxonomy of augment_survey, augmentations can be categorized into model-free, model-based, and policy-based augmentations. Among existing augmentation strategies, we focus on established methods shown to improve robustness or generalization, rather than on the latest or most basic techniques. For the model-free augmentation type, we select AugMix (AugMix9:2021) and PixMix (PixMix:2022), which are augmentations relying on mixing multiple clean images to the original image. For the model-based augmentation type, we choose StyleAug (StyleAug) and DeepAug (DeepAug10:2021), which utilize image-to-image models to diversify the style of the clean images. Finally, among the policy-based augmentation type, we consider RandAugment (RandAugment), which learns the policy for augmenting images.
Distance tinyImageNet ImageNet AugMix DeepAug PixMix RandAug StyleAug AugMix DeepAug PixMix RandAug StyleAug $\gamma=20.0$ 0.0010 0.0008 0.0005 0.0003 0.0000 0.0005 0.0005 0.0002 0.0001 0.0000 $\gamma=30.0$ 0.0054 0.0047 0.0023 0.0014 0.0000 0.0031 0.0028 0.0013 0.0009 0.0000 $\gamma=40.0$ 0.0182 0.0163 0.0084 0.0051 0.0009 0.0113 0.0110 0.0054 0.0034 0.0001 $\gamma=50.0$ 0.0446 0.0399 0.0237 0.0136 0.0087 0.0317 0.0311 0.0169 0.0104 0.0008
Table 4: Empirical CDF $\big(F_{N}(\gamma)\big)$ for selected distance thresholds $(\gamma)$ on (tiny)ImageNet datasets, illustrating sample density near the original image over different augmentation methods. Once again, higher eCDF values mean stronger PSA condition compliance in augmentations. Sorted in descending order.
Tests on Proximal Density
We first measure each augmentation’s proximal density to the original image over different datasets via observing the empirical cumulative distribution function (eCDF) of $L_{2}$ distances on the widely used CIFAR and (tiny)ImageNet datasets (IN; CIFAR). While there exist sophisticated visual-semantic distance metrics (e.g., SSIM, LPIPS), they are not input space metrics. Thus, we have chosen $L_{2}$ distance to reflect our input-parameter space duality. Specifically, we compute how quickly the eCDF accumulates at smaller values, which provides a clear indicator of each augmentation’s density near the original image. Formally, the eCDF of distances $\{d_{i}\}_{i=1}^{N}$ is defined as $F_{N}(\gamma)=\tfrac{1}{N}\sum_{i=1}^{N}$$\mathbf{1}\{d_{i}$$\leq\gamma\}$, where $\mathbf{1}\{\cdot\}$ is the indicator. High values in the low-distance region of $F_{N}(\gamma)$ imply a large fraction of augmented samples remain close to the original image, signaling strong PSA compliance. In contrast, a low accumulation near small $\gamma$ indicates weak adherence. Our analysis reveals that, for the CIFAR datasets (Table 1), AugMix, RandAug, and PixMix fulfill the PSA condition, whereas DeepAug and StyleAug do not comply, ranked in decreasing order of adherence. For (tiny)ImageNet, AugMix, DeepAug, PixMix, and RandAug conform to the PSA condition in descending order, with StyleAug again underperforming.
Tests on Model Flatness
We now explore how an augmentation’s density near the original image affects flatness in model parameters, linking our results from Theorem 1 and 2 with empirical flatness.
Flatness metrics. We assess each model using three quantitative measures: PAC-Bayesian flatness measure $\mu_{\text{PAC-bayes}}$ (PACBayes), local sharpness $\epsilon_{\text{sharp}}$ (eps-sharpness), and Loss-Pass Filter (LPF) metric, a strong indicator of generalization (LPF). Each of these metrics reflects the difference between the proximal loss around a model parameter and its original loss, which aligns well with our theoretical framework. We therefore avoid purely “pinpoint-based” measures (e.g., maximum Hessian eigenvalues or their traces), as they do not fully capture our theorems’ emphasis on local neighborhoods in parameter space. For details, see Appendix C.2.
Experimental results. We train WideResNet-40-2 on CIFAR datasets and ResNet18 on (tiny)ImageNet dataset, following common conventions. For details on training, see Appendix C.2. Table 2 and 5 report the flatness metrics on CIFAR and (tiny)ImageNet. On CIFAR, PSA compliance (Table 1 and 4) is mostly monotone with flatness. Augmentations with larger proximal density (AugMix, RandAug, PixMix) consistently beat ERM, whereas DeepAug and StyleAug give weaker gains. On (tiny)ImageNet, the same trend holds but is less ordered—AugMix and PixMix improve all six metrics, DeepAug and RandAug improve most, and StyleAug again brings little benefit, indicating that PSA is a strong but not exclusive factor that drives flatness.
Tests on Model Robustness
@@
We investigate whether augmentations having rich proximal density (firmly fulfilling PSA condition) fosters robustness against different distribution shifts (Theorem 3).
[…]
Experimental results. We utilize the same WideResNet-40-2 and ResNet18 models used in flatness measures to evaluate their robustness against distribution shifts (Table 3 and 6.) Fundamentally, stronger adherence to the PSA condition reveals an augmentation’s effectiveness in conferring robustness to diverse distribution shifts, consistent with Theorem 3. For further experiments including varying backbone networks (allconvnet; densenet; resnext), refer to Appendix C.4 and D.
Dataset Metrics     ERM    AugMix    DeepAug    PixMix    RandAug    StyleAug tinyImageNet $\mu_{\text{PAC-Bayes}}\downarrow$ 136.83 109.18 (-27.65) 128.41 (-8.42) 106.26 (-30.57) 109.36 (-27.47) 102.72 (-34.11) LPF $\downarrow$ 5.95 3.81 (-2.14) 4.96 (-0.99) 3.40 (-2.55) 4.04 (-1.91) 4.33 (-1.62) $\epsilon_{\text{sharp}}\downarrow$ 15.35 13.09 (-2.26) 25.02 (+9.67) 10.72 (-4.63) 21.30 (+5.95) 35.40 (+20.05) ImageNet $\mu_{\text{PAC-Bayes}}\downarrow$ 219.22 210.63 (-8.59) 211.41 (-7.81) 187.89 (-31.33) 215.58 (-3.64) 235.40 (+16.18) LPF $\downarrow$ 8.70 8.12 (-0.58) 8.67 (-0.03) 8.22 (-0.48) 8.76 (+0.06) 9.27 (+0.57) $\epsilon_{\text{sharp}}\downarrow$ 36.80 30.73 (-6.07) 29.00 (-7.80) 32.76 (-4.04) 36.32 (-0.48) 60.06 (+23.26) Improvement Rate 6/6 5/6 6/6 4/6 2/6
Table 5: Flatness metrics of ERM and different augmentations on (tiny)ImageNet datasets. Overall, augmentations with non-negligible density near the original image (i.e., better satisfying the PSA condition) tend to find flatter minima than ERM. This trend is clear but not perfectly monotone.
   Benchmarks ERM    AugMix    DeepAug    PixMix    RandAug    StyleAug tinyImageNet-C $\downarrow$ 74.76 63.48 (-11.28) 56.14 (-18.62) 61.51 (-13.25) 66.92 (-7.84) 79.42 (+4.66) ImageNet-C $\downarrow$ 69.37 69.34 (-0.03) 56.68 (-12.69) 61.51 (-7.86) 66.88 (-2.49) 75.52 (+6.15) tinyImageNet, $L_{2}$ $\downarrow$ 57.77 61.84 (+4.07) 52.86 (-4.91) 58.38 (+0.61) 63.19 (+5.42) 83.13 (+25.36) tinyImageNet, $L_{\infty}$ $\downarrow$ 99.94 99.93 (-0.01) 99.95 (+0.01) 99.98 (+0.04) 99.98 (+0.04) 100.0 (+0.06) ImageNet, $L_{2}$ $\downarrow$ 76.16 76.06 (-0.10) 75.37 (-0.79) 75.44 (-0.72) 76.09 (-0.07) 87.71 (+11.55) ImageNet, $L_{\infty}$ $\downarrow$ 99.70 99.69 (-0.01) 99.62 (-0.08) 99.63 (-0.07) 99.70 (+0.00) 99.92 (+0.22)     Improvement Rate 5/6 5/6 4/6 3/6 0/6
@@
Table 6: Comparison of various augmentation methods against ERM across multiple robustness scenarios, including common corruptions (tinyImageNet-C, ImageNet-C) and adversarial attacks ($_{2}, LL_{\infty}$). Essentially, the closer an augmentation aligns with the PSA condition, the higher its potential to ensure robustness when confronted with differing distribution shifts.
Discussion and Future Work
@@
  • We first offer a series of theories to explain how augmentation can improve robustness in flat minima viewpoint, end-to-end fashion.
@@
  • We first establish a formal connection between the family of label-invariant augmentations and robustness under general distribution shifts.
@@
    • Previous studies faced limitations such as focusing solely on: i) experimental analysis, ii) particular augmentation, or iii) specific distributional shifts.
@@
    • In contrast, our theories and analyses do not impose such restrictions on augmentations or benchmarks. We validate our findings through experiments across diverse robustness benchmarks.
Despite these contributions, our work has certain limitations. First, we do not provide the kind of extensive experimental validation across many benchmarks as done in technical reports. Second, our theorems do not directly cover label-manipulating augmentations such as Mixup (Mixup:2018) and CutMix (CutMix:2019). Although our framework can be extended to incorporate label-mixing augmentation by setting $\delta$ toward the input mixture, this would require fixing a specific label-mixing rule and would restrict theory generality to the particular augmentation. Thus, we omit label-manipulating augmentations from our main analysis, leaving integration and further schemes to future work.
While not commonly done in augmentation literature, adversarial training (AT) can also be regarded as a specific type of augmentation and serve as a baseline (adv_as_baseline; adv_as_baseline2). AT improves robustness to norm-bounded attacks but stays brittle under large distribution shifts (e.g., common corruptions). We hypothesize that such a phenomenon occurs since AT has limitations in promoting large $b$-flat minima due to diminutive perturbations. Therefore, combining AT with augmentations that generate samples in both short-distance and long-distance range (e.g., AugMix) helps counteract its vulnerability to large shifts, by effectively boosting $\gamma_{A}\gg 0$ (Table 78).
Dataset Metrics    AT    AT + AugMix CIFAR-10 $\mu_{\text{PAC-Bayes}}\downarrow$    158.15 113.78 (-44.37) LPF $\downarrow$    1.25 0.58 (-0.67) $\epsilon_{\text{sharp}}$ $\downarrow$    16.36 10.29 (-6.07) CIFAR-100 $\mu_{\text{PAC-Bayes}}\downarrow$    207.90 161.89 (-46.01) LPF $\downarrow$    2.95 2.11 (-0.84) $\epsilon_{\text{sharp}}\downarrow$    20.81 10.34 (-10.47)
Table 7: Flatness evaluations of AT (FGSM, $\epsilon=8/255$) and AT + AugMix on CIFAR-10/100.
Robustness Benchmarks AT AT + AugMix CIFAR-10-C $\downarrow$ 24.12 15.39 (-8.73) Common CIFAR-10-$\overline{\text{C}}$ $\downarrow$ 3.25 1.30 (-1.95) Corruption CIFAR-100-C $\downarrow$ 52.50 44.03 (-8.47) CIFAR-100-$\overline{\text{C}}$ $\downarrow$ 7.36 3.74 (-3.62) CIFAR-10, $L_{2}$ $\downarrow$ 69.24 62.86 (-6.38) Adversarial CIFAR-10, $L_{\infty}$ $\downarrow$ 94.55 98.40 (+3.85) Attacks CIFAR-100, $L_{2}$ $\downarrow$ 89.28 88.60 (-0.68) CIFAR-100, $L_{\infty}$ $\downarrow$ 98.42 99.75 (+1.33)
Finally, it is well known that models trained with flat-minima optimizers, such as Sharpness-Aware Minimization (SAM) (SAM:2021) and Stochastic Weight Averaging (SWA) (SWA:2018), generalize better under distribution shifts. However, the mechanism by which data augmentation fosters flat minima remains underexplored. We bridge this gap by introducing a simple sufficient ”proximal support” condition and proving that PSA encourages flat minima. See Appendix D for further discussions on applications, flat-minima optimizers, and adversarial training.
Related Works
Data Augmentations
Existing augmentation techniques can be categorized into model-free, model-based, and policy-based approaches, following the classification proposed in augment_survey.
Model-free augmentations are further subdivided into single-image and multi-image techniques. Single-image methods encompass basic image operations like translation, rotation, and color jitter, as well as masking strategies such as CutOut (CutOut) and Hide-and-Seek (HideNSeek). These transformations have been shown to enhance model performance on target data distributions efficiently and with low overhead. Multi-image augmentations involve composing multiple operations and blending elements from different images. Among label-manipulating multi-image methods are techniques that combine pairs of distinct images and their labels, including Mixup (Mixup:2018) and CutMix (CutMix:2019). These approaches are widely recognized for their ability to boost model performance. Subsequently, label-preserving multi-image methods like AugMix (AugMix9:2021) and PixMix (PixMix:2022) have demonstrated superior effectiveness in enhancing model robustness.
Model-based augmentations leverage pretrained models to produce augmented data. Several techniques employing generative models, such as CGAN (CGAN) and its variants (CGAN_var1; CGAN_var2; CGAN_var3; CGAN_var4), are designed to mitigate data imbalance issues. Additional methods, including DeepAugment (DeepAug10:2021), ANT (ANT11:2020), and StyleAug (StyleAug), focus on improving classifier robustness to common corruptions, adversarial attacks, or domain shifts. More recently, pretrained diffusion models combined with effective prompt engineering (Diffusion1; Diffusion2) has proven successful in enhancing performance across various vision tasks.
Policy-based augmentations focus on designing an automatic way to determine the optimal augmentation strategies by employing reinforcement learning or adversarial training. From a pioneering method, AutoAugment (AutoAugment) that utilizes reinforcement learning for finding the best augmentation strategies, subsequent works such as Fast AA (FastAA), Faster AA (FasterAA), and RandAugment (RandAugment) aim to enhance both the efficiency of policy search and the model performance. Adversarial training-based augmentation strategies, including AdaTransform (AdaTransform), Adversarial AA (AdvAA), and AugMax (AugMax) leverage adversarial perturbations which maximally disturbs samples to be misclassified into other labels, finally leading to improve model robustness against unseen domains.
Augmentations and Model Robustness
aug_analysis_cc1 has provided a game-theoretic perspective on augmentations and introduces a new proxy metric for measuring common corruption robustness. ME-ADA and ADA have theoretically found out that adversarial perturbations in the latent space can simulate worst-case distributional shifts in the data. AugmentNAdvGood have empirically found out that Mixup and CutOut with model weight averaging is shown to improve adversarial robustness. boundary_thickness has shown that Mixup enlarges boundary thickness, the marginal space between differently labeled samples, which is deemed to be highly correlated with adversarial robustness and common corruptions. AugFourier interpret the augmentation-originated gains in model robustness by explaining that augmentations make deep models utilize both high and low frequency information of images so as to enhance model robustness against data corruptions. NoLabelNAdvGood1 and NoLabelNAdvGood2 have theoretically shown that utilizing unlabeled data in training can improve adversarial robustness. DeepAug10:2021 have empirically found out that exploiting diverse augmentations together improves robustness against both adversarial and common corruptions.
Nonetheless, most of the prior interpretations of how augmentation contributes to model robustness are either confined to specific types of augmentations and robustness or empirical analysis. None of the prior works generally explain how augmentations can theoretically improve model robustness across diverse distributional shifts.
Flat Minima and Robustness
The relationship between flat minima in the loss landscape and improved generalization performance has been well-established in several seminal works (PAC_DG1; SWAD:2021; zhang2024duality; CL_flat; FL_flat). Flat minima refer to regions in the optimization landscape where the loss function remains relatively stable and insensitive to small perturbations in the model parameters, which often correlate with better generalization to unseen data. For instance, PAC_DG1 establishes a rigorous theoretical link between these flat minima and superior generalization capabilities in over-parameterized neural networks, achieving this by refining the PAC-Bayes generalization bound originally proposed in PAC_DG2. This refinement provides a more precise quantification of how flatter regions can lead to models that perform more reliably on out-of-distribution data.
In a similar vein, SWAD:2021 demonstrates that flat minima contribute to tighter bounds on generalization error, supported by both theoretical analysis and practical experiments. Their approach highlights dense weight averaging as a practical and efficient heuristic for locating such minima during training, showing through empirical validation that it consistently enhances model performance across various benchmarks. Furthermore, zhang2024duality investigates how sharpness-aware minimization (SAM) relates to adversarial training (AT), demonstrating that SAM can replicate AT’s behavior under the restrictive setting of Gaussian inputs and linear models. Because their theoretical analysis is confined to binary classification and Gaussian data, our theorems and the PSA condition have an edge in applying to general data distributions.
Despite these advances, prior work still leaves the open question of how data augmentation itself steers optimization toward flat minima. Our contribution addresses this critical gap by forging a direct connection between data augmentation strategies and the identification of flat minima, while also demonstrating their pivotal role in bolstering model robustness against a wide array of general distribution shifts, such as those encountered in real-world scenarios involving domain variations or noisy inputs.
Other Related Works
jiang2023chasing investigates how to preserve fairness under distribution shifts by reinterpreting such shifts as equivalent perturbations to both model weights and inputs, employing a transportation function. Still, their work focuses on group fairness with respect to demographic parity and equal opportunity, and demonstrates only the feasibility of this equivalence without further extensions.
flat_minima_cc shows in binary classification that flatter minima mitigate the boundary tilting problem (boundary_tilting), and empirically finds these minima correlate with greater robustness to common corruptions. However, the theory remains limited, as the boundary tilting problem is closely tied to adversarial robustness and is not explored beyond binary classification.
Conclusion
Our framework connects data augmentation to robustness through flat minima viewpoint under data shifts. Augmentations meeting the PSA condition form broad, flat regions in parameter space, while others provide limited protection. Experiments on CIFAR and ImageNet substantiate our theorems, and we expect our work to inspire next-generation augmentation methods, including foundation model-based.
Acknowledgments